Solve the system:
5x + 4y - 3z = 15
x - 5z = 11
2x + 3z = -4

Answer by stanbon(48545)

Solve the system:
5x + 4y - 3z = 15
x - 5z = 11
2x + 3z = -4
--------------
Form the 3x3 coefficient matrix = A
Form the 3x1 constant matrix = B
--------------
EQUATION:
A^-1B = [1,1,-2]
x=1,y=1,z=-2
=============
Cheers,
Stan H.